in 1 hour 3/4 pint of chemicals can be filtered. How long in hours will it take to filter 2 1/2 pints?
Answer by mananth(16949)

in 1 hour 3/4 pint of chemicals can be filtered. How long in hours will it take to filter 2 1/2 pints?
3/4 pints fitered in 1 hour
2 1/2 pints in ?
2 1/2 = 5 /2 pints
5/2 *1 / 3/4
5/2 * 4/3
= 10 / 3 hours or 3 1/3 hours